OpenStudy (anonymous):

Please help me! h^2+4=2h solve for h

OpenStudy (anonymous):

\[h^2-2h+4=0\] you can use the quadratic formula, but it is easier to complete the square for this one

OpenStudy (anonymous):

h=-2 i think

OpenStudy (anonymous):

yes, but how to solve :)

OpenStudy (anonymous):

\[h^2-2h+4=0\]\[h^2-2h=-4\]\[(h-1)^2=-4+1=-3\]\[h-1=\pm\sqrt{-3}=\pm\sqrt{3i}\]\[h=1\pm\sqrt{3}i\]

OpenStudy (anonymous):

D < 0 the roots not real

OpenStudy (anonymous):

solutions are complex numbers so if you are working with real numbers you would say "no solution" but if you allow complex solutions the answer and method is above

OpenStudy (anonymous):

is 1 necessary infront of +/_square root of 3 i ?

OpenStudy (anonymous):

you can write \(\sqrt{3}i+1\) or \(-\sqrt{3}i+1\) if you like but standard form of a complex number is \(a+bi\)

OpenStudy (anonymous):

you put the real part first and the "imaginary part" second, so yes you should probably write \(1+\sqrt{3}i\) and \(1-\sqrt{3}i\)

OpenStudy (anonymous):

Thank you so much! I really appreciate that!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

that satellite person sounds right